OpenStudy (anonymous):

Write an equation that could be used to find the value of b. Answer 382 = 22 + 32 – 2(2)(3)(cos b°) b2 = 22 + 32 – 2(2)(3)(cos 38°) 22 = b2 + 32 – 2(b)(3)(cos 38°) 32 = 22 + b2 – 2(2)(b)(cos 38°)

OpenStudy (anonymous):

ummm..we would appreciate if there is a picture of some sort.

OpenStudy (anonymous):

alright :) hold on one sec

OpenStudy (anonymous):

@dg98 :)

OpenStudy (anonymous):

oh that's easy! ok so cosine law

OpenStudy (anonymous):

a^2=b^2+c^2-2bc cosA

OpenStudy (anonymous):

or for this question...b^2=c^2+a^2-2ac cosB

OpenStudy (anonymous):

=9+4-12cos38

OpenStudy (anonymous):

oops i was solving it. b^2=3^2+2^2-2(3)(2)cos38

OpenStudy (anonymous):

thanks!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

you're welcome! :)